# Protecting Human Activity Signatures in Compressed IEEE 802.11 CSI Feedback

### v1

@@ IV Numerical Simulations
We evaluate the impact of DP Givens–angle quantization scheme on both downlink beamforming utility and the adversary’s ability to estimate user motion speed from the reported angular parameters. The downlink channel $\mathbf{H}(t)$ follows a standard time-varying multipath fading model at each coherence block, the receiver computes the right-singular vector of $\mathbf{H}(t)$ and encodes it into Givens rotation and phase angles following the 802.11 compressed beamforming format. We apply our DP stochastic quantizer to the true angles, drawing each reported angle $\widehat{\psi}$ or $\widehat{\phi}$ from an $\varepsilon$-DP distribution centered at its true value. The AP uses the privatized angles to reconstruct a semi-unitary transmit beamformer $\widehat{\mathbf{V}}$, while a passive adversary receives the same angles and attempts speed estimation following the micro-Doppler extraction pipeline in [cominelli2024physical].
Impact on Bit Error Rate (BER). In Fig. 5, we report the impact of DP–SQ on link performance as measured by BER. The first row corresponds to low-resolution (1-bit) CSI feedback, while the second row shows high-resolution (3-bit) feedback. As expected, higher modulation orders or lower angular resolutions increase the system’s sensitivity to DP-induced beam misalignment, resulting in elevated BER across SNR levels. Notably, in the high-resolution setting, the proposed DP–SQ mechanism achieves BER performance nearly indistinguishable from the non-private baseline, demonstrating that strong privacy can be attained with minimal loss when sufficient angular resolution is available.
In Fig. 6, we visualize the received constellation points to provide intuition behind these trends. SVD beamforming produces tightly clustered decision points around the ideal 16-QAM constellation. Deterministic quantization introduces moderate angular distortion, broadening the symbol clouds. The DP–SQ mechanism injects controlled stochasticity into the angular representation, further dispersing the received symbols while offering formal DP guarantees.
@@
Next, in Fig. 7, we report the average beamforming gain $G$ over $5000$ Rayleigh fading realizations for several IEEE 802.11 antenna configurations with $(N_{t},N_{r})\in{(2,1),(2,2),(2,3),(2,4),(2,8)}$ and a single spatial stream ($N_{s}=1$). Fig. 7 reports the mean gain achieved by three schemes: (i) deterministic midpoint quantization of the Givens angles, (ii) the proposed differentially private stochastic quantization (DP–SQ), and (iii) BeamDancer-style randomized beamforming [cominelli2024physical]. The perfect SVD beamformer provides the reference gain $G=1$. Deterministic quantization incurs moderate loss due to finite angular resolution. The proposed DP–SQ mechanism introduces controlled stochastic perturbations governed by $(\varepsilon_{\phi},\varepsilon_{\psi})$, producing a tunable privacy–utility tradeoff: smaller values of $\varepsilon$ increase privacy but yield additional beam misalignment and corresponding gain reduction. BeamDancer exhibits the largest degradation, as its high-variance integer-bin perturbations cause substantial angular distortion. Across all antenna configurations, the trend is consistent: privacy-driven angular randomization directly impacts achievable array gain, with the effect magnified in larger arrays—yet the proposed DP–SQ scheme preserves compatibility with the 802.11 feedback format while offering significantly improved privacy–utility behavior compared to prior randomized approaches.
@@
Robustness to Adversarial Estimation. Fig. 8 demonstrates the effectiveness of the proposed privacy-based perturbations over time: despite the injected noise on the compressed CSI, the normalized per-stream beamforming gain generally remains close. Fig. 9 further quantifies this, where the median gain obtained from the distorted CSI is $0.89$ (compared to the deterministic median gain of $0.97$ and never drops below $0.52$. This indicates that the DP mechanism preserves most of the communication utility. At the same time, the same perturbations substantially degrade the eavesdropper’s ability to track the user’s motion, leading to visibly biased and inaccurate velocity trajectories compared to the ground truth. This illustrates the key privacy–utility tradeoff achieved by our design: the beamformer retains near-optimal precoding performance from the CSI privatized feedback, while an adversary observing the perturbed CSI suffers a pronounced loss in speed-estimation accuracy from the reduced information leakage.
@@
It is worth emphasizing that the privacy guarantee of the proposed mechanism is per-shot: each CSI report is individually protected under $\varepsilon$-DP. In our setup, the adversary obtains one CSI measurement every $10^{-3}$ seconds over a total communication duration of $T=10$ seconds, resulting in 10,00010{,}000 DP-protected releases. By applying the advanced composition theorem of DP [dwork2014algorithmic], the cumulative privacy loss grows sublinearly, yielding an overall leakage on the order of $\varepsilon_{\mathrm{total}}={O}(\sqrt{T}\,\cdot\max\{\varepsilon_{\psi},\varepsilon_{\phi}\})$, demonstrating that even over long communication intervals, the aggregated DP leakage remains well controlled.

### v2

@@ III-C DP Stochastic Quantization (DP-SQ) Mechanism
Remark 1.
The quantities $\sigma_{\psi}^{2}$ and $\sigma_{\phi}^{2}$ correspond to the per-angle mean squared errors induced by the DP stochastic quantization mechanism, thereby explicitly linking the privacy parameter $\epsilon$ to the resulting beamforming distortion.
We next show that the proposed DP-SQ mechanism preserves the orthonormality structure of the SVD-based beamforming matrix.
Orthonormality under DP-SQ Mechanism. To make the argument concrete, consider the single-stream case with $N_{t}=2$ and $N_{r}=1$. In this setting, any unit-norm beamforming vector admits the angular parametrization
$\mathbf{v}(\psi,\phi)=\begin{bmatrix}\cos\psi\\ e^{j\phi}\sin\psi\end{bmatrix},\qquad\psi\in[0,\tfrac{\pi}{2}],\ \phi\in[-\pi,\pi).$ (18)
[…]
$\|\widehat{\mathbf{v}}\|_{2}^{2}=|\cos\widehat{\psi}|^{2}+|e^{j\widehat{\phi}}\sin\widehat{\psi}|^{2}=\cos^{2}\widehat{\psi}+\sin^{2}\widehat{\psi}=1,$
where the equality follows from the trigonometric identity $\cos^{2}x+\sin^{2}x=1$ and the unit-modulus property $|e^{j\widehat{\phi}}|=1$ for all $\widehat{\phi}\in[-\pi,\pi)$. Thus, stochastic quantization of the angles preserves the unit-norm (and hence orthonormality) constraint exactly, while affecting only the orientation of the beamforming direction. The same reasoning extends to higher dimensions, where the beamforming matrix is constructed as a product of Givens rotations and phase factors, each of which remains unitary for all quantized angle values.
III-D Performance Guarantees
We now characterize the bit error rate (BER) and achievable rate as a function of the proposed privacy-preserving perturbation applied to the beamforming matrix. The analysis proceeds by first mapping the resulting subspace distortion into an effective SNR, which directly governs both performance metrics. We start with the single stream case $N_{s}=1$. The STA reconstructs a unit-norm beamforming vector $\widehat{\mathbf{v}}\in\mathbb{C}^{N_{t}}$ from the compressed CSI feedback (under the DP-SQ mechanism), and the received SNR is
[…]
$=\frac{P}{N_{0}}\,\cdot\widehat{\mathbf{v}}^{H}\mathbf{H}^{H}\mathbf{H}\,\widehat{\mathbf{v}}$
$=\frac{P}{N_{0}}\,\cdot\widehat{\mathbf{v}}^{H}\mathbf{V}\bm{\Sigma}^{2}\mathbf{V}^{H}\widehat{\mathbf{v}}$ (19)
$=\frac{P}{N_{0}}\cdot\sum_{i\geq 1}\sigma_{i}^{2}\,\cdot\big|\langle\mathbf{v}_{i},\widehat{\mathbf{v}}\rangle\big|^{2},$ (20)
where (19) follows from $\mathbf{H}^{H}\mathbf{H}=\mathbf{V}\bm{\Sigma}^{2}\mathbf{V}^{H}$, and (20) follows by expanding $\widehat{\mathbf{v}}$ in the orthonormal basis $\{\mathbf{v}_{i}\}$. Now, let $\mathbf{v}^{\star}\triangleq\mathbf{v}_{1}$ denote the dominant right-singular vector. Separating the dominant contribution yields
$\gamma(\widehat{\mathbf{v}})=\frac{P}{N_{0}}\cdot\left(\sigma_{1}^{2}\,\cdot\big|\langle\mathbf{v}^{\star},\widehat{\mathbf{v}}\rangle\big|^{2}+\sum_{i\geq 2}\sigma_{i}^{2}\,\cdot\big|\langle\mathbf{v}_{i},\widehat{\mathbf{v}}\rangle\big|^{2}\right).$ (21)
Since $\widehat{\mathbf{v}}$ is unit norm and $\{\mathbf{v}_{i}\}$ is orthonormal, we have
$\sum_{i\geq 1}\big|\langle\mathbf{v}_{i},\widehat{\mathbf{v}}\rangle\big|^{2}=1,\qquad\sum_{i\geq 2}\big|\langle\mathbf{v}_{i},\widehat{\mathbf{v}}\rangle\big|^{2}=1-\big|\langle\mathbf{v}^{\star},\widehat{\mathbf{v}}\rangle\big|^{2}.$ (22)
For rank-one beamforming, the squared chordal distance satisfies
$d_{\mathrm{chord}}^{2}(\mathbf{v}^{\star},\widehat{\mathbf{v}})=1-\big|\langle\mathbf{v}^{\star},\widehat{\mathbf{v}}\rangle\big|^{2}.$ (23)
Using the fact that $\sigma_{i}^{2}\geq\sigma_{2}^{2}$ for all $i\geq 2$ and (22), we further get
[…]
To this end, we obtain the following lower bound:
[…]
where
$\gamma^{\star}(\mathbf{H})\triangleq\frac{P}{N_{0}}\cdot\sigma_{1}^{2}(\mathbf{H}),\qquad\gamma_{2}(\mathbf{H})\triangleq\frac{P}{N_{0}}\cdot\sigma_{2}^{2}(\mathbf{H}).$
In particular, since $\gamma_{2}(\mathbf{H})\geq 0$, a simpler (but slightly looser) bound is obtained by discarding the second term:
$\gamma(\widehat{\mathbf{v}})\ \geq\ \gamma^{\star}(\mathbf{H})\bigl(1-d_{\mathrm{chord}}^{2}(\mathbf{v}^{\star},\widehat{\mathbf{v}})\bigr)\ \triangleq\ \gamma_{\mathrm{LB}}(\mathbf{H}).$ (24)
III-D1 Ergodic Achievable Rate
We now characterize the achievable rate degradation induced solely by the proposed DP-SQ mechanism. The instantaneous achievable rate (in bits-per-second-per Hz) is given by
[…]
Since the main source of randomness we are interested in, within a coherence block is the DP-SQ mechanism, we define
[…]
This corresponds to the achievable rate averaged over the stochastic feedback mechanism for a fixed channel realization. If $\mathbf{H}$ varies ergodically across blocks, an additional expectation over $\mathbf{H}$ yields the classical ergodic rate [4].
From Section IV-A, the received SNR satisfies the affine lower bound
$\gamma(\widehat{\mathbf{v}})\ \geq\ \gamma^{\star}(1-d)+\gamma_{2}d,$ (26)
where $d\triangleq d_{\mathrm{chord}}^{2}(\mathbf{v}^{\star},\widehat{\mathbf{v}})\in[0,1]$. While (26) is tight, the resulting rate function is concave in $d$, which prevents a direct Jensen-type lower bound on $\bar{R}$.
Smooth SNR lower bound. To enable tractable averaging, we introduce the conservative bound
$\gamma(\widehat{\mathbf{v}})\ \geq\ \frac{\gamma^{\star}}{1+\eta d},$ (27)
where $\eta>0$ is chosen such that $\eta\geq\frac{\gamma^{\star}-\gamma_{2}}{\gamma^{\star}}.$ This form is tight at $d=0$ and upper-bounds the worst-case SNR degradation induced by subspace mis-alignment. We next present a lower bound guarantee on the achievable rate.
Theorem 2 (Ergodic rate lower bound under DP-SQ).
[…]
$\bar{R}\ \geq\ \log_{2}\!\left(1+\frac{\gamma^{\star}}{1+\eta\,\mathbb{E}_{\mathrm{DP}}[d]}\right).$ (28)
Proof.
Substituting (27) into (25) yields
$\bar{R}\ \geq\ \mathbb{E}_{\mathrm{DP}}\!\left[\log_{2}\!\left(1+\frac{\gamma^{\star}}{1+\eta d}\right)\right].$
The function $g(d)=\log_{2}\!\left(1+\frac{\gamma^{\star}}{1+\eta d}\right)$ is convex for $d\geq 0$. Applying Jensen’s inequality completes the proof. ∎
From Theorem 1, the DP-SQ mechanism ensures $\mathbb{E}_{\mathrm{DP}}[d]\leq\mathsf{D}(\epsilon_{\psi},\epsilon_{\phi}).$ Substituting into (28) yields
$\bar{R}\ \geq\ \log_{2}\!\left(1+\frac{\gamma^{\star}}{1+\eta\,\mathsf{D}(\epsilon_{\psi},\epsilon_{\phi})}\right).$ (29)
The above lower bound establishes an explicit privacy–rate tradeoff: tighter privacy budgets increase the expected subspace distortion $\mathbb{E}[d]$, which reduces the effective SNR through the factor $(1+\eta d)^{-1}$ and consequently degrades the achievable rate. Unlike the BER metric, this degradation is logarithmic in SNR, implying that rate is more robust to moderate privacy-induced perturbations, especially in the high-SNR regime.
III-D2 Bit Error Rate
For square $M$-QAM with Gray mapping over an AWGN channel, the instantaneous bit error rate (BER) admits the standard approximation [6]:
$P_{b}(\widehat{\mathbf{v}})\ \approx\ c_{M}\,\cdot Q\!\Bigg(\sqrt{\frac{3\,\gamma(\widehat{\mathbf{v}})}{M-1}}\Bigg),\quad c_{M}\triangleq\frac{4}{\log_{2}M}\Big(1-\frac{1}{\sqrt{M}}\Big),$ (30)
We next show the impact of the DP-SQ mechanism on the BER.
Lemma 2 (BER bound under DP-SQ).
[…]
$P_{b}(\widehat{\mathbf{v}})\ \leq\ c_{M}\,\cdot Q\!\Bigg(\sqrt{\frac{3\,\gamma^{\star}(1-d)}{M-1}}\Bigg).$ (31)
Proof.
The result follows by substituting the SNR lower bound into the monotone decreasing function (30). ∎
Average BER under DP-SQ randomness. Since the DP-SQ mechanism induces randomness in $\widehat{\mathbf{v}}$, the relevant performance metric is the average BER
[…]
Using Lemma 2, we obtain the bound
$\bar{P}_{b}\ \leq\ c_{M}\,\cdot\mathbb{E}\!\left[Q\!\Bigg(\sqrt{\frac{3\,\gamma^{\star}(1-d)}{M-1}}\Bigg)\right],$ (33)
where the expectation is taken over the distribution of the chordal distortion $d$ induced by DP-SQ. Using the convexity of $Q(\sqrt{x})$ for $x>0$ and Jensen’s inequality, we further obtain
$\bar{P}_{b}\ \leq\ c_{M}\,\cdot Q\!\Bigg(\sqrt{\frac{3\,\gamma^{\star}(1-\mathbb{E}[d])}{M-1}}\Bigg).$ (34)
Equations (31)-(34) establish a direct and interpretable link between privacy and reliability. The DP-SQ mechanism increases the expected subspace distortion $\mathbb{E}[d]$, which reduces the effective SNR through a multiplicative factor (1d)(1-d) and results in an exponential degradation in BER via the $Q$-function. This characterization is particularly sharp for high-order constellations, where reliability is highly sensitive to SNR loss.
III-E Global DP via Full-Support Randomization
The DP-SQ mechanism introduced above provides local indistinguishability within each quantization interval, but does not satisfy global DP (c.f. Definition 1) over the entire angular domain due to its two-point output support. In particular, inputs from different quantization cells may induce output distributions with disjoint supports.
To address this limitation, we introduce a globally private extension that assigns non-zero probability to all reconstruction levels while preserving locality around the true angle. We refer to this globally private extension as DP-GSQ, as it combines stochastic quantization with a full-support geometric randomization kernel.
DP-GSQ Mechanism. Let $\mathcal{Q}=\{q_{1},\dots,q_{L}\}$ denote the uniform quantization grid. For a given input angle $a$, let $q_{i}\leq a\leq q_{i+1}$ denote the bracketing levels. Define the stochastic interpolation weights
$\lambda_{i}(a)=\frac{q_{i+1}-a}{q_{i+1}-q_{i}},\qquad\lambda_{i+1}(a)=\frac{a-q_{i}}{q_{i+1}-q_{i}}.$
We define a truncated geometric kernel centered at level $q_{j}$ as follows:
$G_{\tau}(k\mid j)=\frac{(1-\tau)\,\tau^{d(k,j)}}{Z_{j}(\tau)},\qquad k\in\{1,\dots,L\},$
where $0<\tau<1$, $d(k,j)$ is the index distance (or circular distance for $\phi$), and $Z_{j}(\tau)$ is a normalization constant. The globally private stochastic quantizer releases $\widehat{a}\in\mathcal{Q}$ according to
$\Pr(\widehat{a}=q_{k}\mid a)=\lambda_{i}(a)\,G_{\tau}(k\mid i)+\lambda_{i+1}(a)\,G_{\tau}(k\mid i+1).$ (35)
Fig. 6 illustrates the key distinction between the two mechanisms. While local DP-SQ perturbs only within the bracketing quantization cell, DP-GSQ ($\tau=0.35$) spreads a geometrically decaying tail over the entire angle codebook. This full-support randomization is precisely what enables a finite global DP guarantee.
Figure 6: Visualization of deterministic quantization, local DP-SQ, and the proposed globally private DP-GSQ mechanism for a fixed input angle. (a) Deterministic quantization places all probability mass on a single reconstruction level. (b) Local DP-SQ randomizes only between the two neighboring levels. (c) DP-GSQ uses truncated geometric kernels centered at the two bracketing levels. (d) The final DP-GSQ output distribution is a convex combination of these kernels and assigns non-zero probability to all quantization levels. This full-support property eliminates zero-probability outputs and enables a global differential privacy guarantee over the entire angular domain.
Since $G_{\tau}(k\mid j)>0$ for all $k,j$, every output level has non-zero probability under any input. Therefore, the mechanism satisfies $\varepsilon$-local DP over the full angular domain.
Theorem 3 (Global DP guarantee).
The mechanism in (35) satisfies $\varepsilon_{a}$-local DP, i.e.,
$\Pr(\widehat{a}\in S\mid a)\leq e^{\varepsilon_{a}}\Pr(\widehat{a}\in S\mid a^{\prime}),\quad\forall a,a^{\prime},\ \forall S\subseteq\mathcal{Q},$
where
$\varepsilon_{a}\leq\log\left(\frac{\max_{j,k}G_{\tau}(k\mid j)}{\min_{j,k}G_{\tau}(k\mid j)}\right).$
Proof Sketch. The proof follows by bounding the likelihood ratio
$\sup_{a,a^{\prime}}\sup_{k}\frac{\Pr(\widehat{a}=q_{k}\mid a)}{\Pr(\widehat{a}=q_{k}\mid a^{\prime})}.$
Since $\Pr(\widehat{a}=q_{k}\mid a)$ is a convex combination of strictly positive kernels $G_{\tau}(k\mid i)$ and $G_{\tau}(k\mid i+1)$, the ratio is bounded by the ratio between the maximum and minimum values of $G_{\tau}(k\mid j)$ across all indices. This yields the stated bound.
III-E1 Operational interpretation of DP-GSQ
The kernel parameter $\tau$ controls the concentration of the output distribution over the quantization alphabet. In particular, the truncated geometric kernel assigns probability mass according to $G_{\tau}(k\mid j)\propto\tau^{d(k,j)}$, so that $\tau$ acts as a density-flattening parameter. In the regime $\tau\to 0$, the kernel becomes sharply concentrated around the bracketing reconstruction levels, and the DP-GSQ mechanism behaves similarly to a localized stochastic quantizer with negligible tails. In contrast, as $\tau\to 1$, the geometric decay vanishes and the output distribution approaches a uniform distribution over all quantization levels. Consequently, DP-GSQ continuously interpolates between localized stochastic quantization and full-support randomization, providing a tunable mechanism for controlling the privacy–utility tradeoff.
As $\tau$ increases, the output distribution becomes more uniform, reducing the distinguishability between different inputs and thereby decreasing $\varepsilon_{a}(\tau)$, $a\in\{\theta,\phi\}$. Conversely, smaller values of $\tau$ concentrate probability mass near the true angle, increasing distinguishability and hence the privacy leakage. This establishes a direct and interpretable relationship between the kernel parameter $\tau$, the induced privacy level, and the resulting distortion.
Distortion analysis. Under the proposed mechanism, the per-angle distortion is given by
$\mathbb{E}\!\left[(\widehat{a}-a)^{2}\mid a\right]=\sum_{k=1}^{L}(q_{k}-a)^{2}\,\Pr(\widehat{a}=q_{k}\mid a).$
Averaging over $a$ within each quantization interval yields the effective angular distortion
$\sigma_{a,G}^{2}=\frac{1}{\Delta}\int_{q_{i}}^{q_{i+1}}\sum_{k=1}^{L}(q_{k}-a)^{2}\,\Pr(\widehat{a}=q_{k}\mid a)\,da.$
Remark 2.
Compared to the original DP-SQ mechanism, the globally private extension ensures rigorous DP over the entire angular domain at the cost of increased distortion, as probability mass is spread across all quantization levels. This highlights a fundamental tradeoff between global privacy guarantees and beamforming accuracy.
Proposition 1 (MSE bound for DP–GSQ).
Consider the globally private stochastic quantization mechanism in (35) over a uniform grid $\mathcal{Q}=\{q_{1},\dots,q_{L}\}$ with spacing $\Delta$. Define
$M_{\tau}(j)\triangleq\sum_{k=1}^{L}(q_{k}-q_{j})^{2}G_{\tau}(k\mid j).$
Then, for any input $a\in[q_{i},q_{i+1}]$, the conditional distortion satisfies
$\mathbb{E}\!\left[(\widehat{a}-a)^{2}\mid a\right]\leq 2\max_{1\leq j\leq L}M_{\tau}(j)+\frac{\Delta^{2}}{2}.$
Moreover, if $a$ is uniformly distributed within its quantization interval, then
$\mathbb{E}\!\left[(\widehat{a}-a)^{2}\right]\leq 2\max_{1\leq j\leq L}M_{\tau}(j)+\frac{\Delta^{2}}{6}.$
Proof.
By definition,
$\Pr(\widehat{a}=q_{k}\mid a)=\lambda_{i}(a)G_{\tau}(k\mid i)+\lambda_{i+1}(a)G_{\tau}(k\mid i+1).$
Hence
$\mathbb{E}[(\widehat{a}-a)^{2}\mid a]=\lambda_{i}(a)\sum_{k}(q_{k}-a)^{2}G_{\tau}(k\mid i)$
$+\lambda_{i+1}(a)\sum_{k}(q_{k}-a)^{2}G_{\tau}(k\mid i+1).$
Using
$(q_{k}-a)^{2}\leq 2(q_{k}-q_{j})^{2}+2(q_{j}-a)^{2},$
for $j\in\{i,i+1\}$, we obtain
$\sum_{k}(q_{k}-a)^{2}G_{\tau}(k\mid j)\leq 2M_{\tau}(j)+2(q_{j}-a)^{2}.$
Substituting this into the previous expression gives
$\mathbb{E}[(\widehat{a}-a)^{2}\mid a]\leq 2\max_{j}M_{\tau}(j)$
$+2\bigl(\lambda_{i}(a)(a-q_{i})^{2}+\lambda_{i+1}(a)(q_{i+1}-a)^{2}\bigr).$ (36)
Since the last term is upper bounded by $\Delta^{2}/2$, the conditional bound follows. Averaging over $a$ uniformly within the interval yields the second statement. ∎
Corollary 1 (Subspace distortion under DP-GSQ).
Let $\sigma_{\psi,\mathrm{G}}^{2}\triangleq\mathbb{E}[(\widehat{\psi}-\psi)^{2}]$ and $\sigma_{\phi,\mathrm{G}}^{2}\triangleq\mathbb{E}[(\widehat{\phi}-\phi)^{2}]$ denote the per-angle mean squared errors induced by the DP-GSQ mechanism. Then the expected squared chordal distance between the ideal subspace $\mathcal{S}^{\star}$ and the reconstructed subspace $\widehat{\mathcal{S}}_{\mathrm{G}}$ satisfies
$\mathbb{E}\!\left[d_{\mathrm{chord}}^{2}(\mathcal{S}^{\star},\widehat{\mathcal{S}}_{\mathrm{G}})\right]\leq d_{q}^{2}+2\cdot N_{s}\cdot N_{\mathrm{tot}}\cdot\bigl(\sigma_{\psi,\mathrm{G}}^{2}+\sigma_{\phi,\mathrm{G}}^{2}\bigr).$
The result follows by applying the same Givens-cascade perturbation argument as in Theorem 1, with the per-angle error variances replaced by those induced by the DP-GSQ mechanism. The corresponding achievable rate and BER characterizations follow along the same lines.
@@ IV Numerical Simulations
In this section, we evaluate the impact of DP Givens angle quantization scheme on both downlink beamforming utility and the adversary’s ability to estimate user motion speed from the reported CSI angular parameters. The downlink channel $\mathbf{H}(t)$ follows a standard time-varying multipath fading model at each coherence block, the receiver computes the right-singular vector of $\mathbf{H}(t)$ and encodes it into Givens rotation and phase angles following the 802.11 compressed beamforming format. We apply our DP stochastic quantizer to the true angles, drawing each reported angle $\widehat{\psi}$ or $\widehat{\phi}$ from an $\varepsilon$-DP distribution centered at its true value. The AP uses the privatized angles to reconstruct a semi-unitary transmit beamformer $\widehat{\mathbf{V}}$, while a passive adversary receives the same angles and attempts speed estimation following the micro-Doppler extraction pipeline in [2].
Impact on Bit Error Rate (BER). In Fig. 7, we report the impact of DP–SQ on link performance as measured by BER. The first row corresponds to low-resolution (1-bit) CSI feedback, while the second row shows high-resolution (3-bit) feedback. As expected, higher modulation orders or lower angular resolutions increase the system’s sensitivity to DP-induced beam mis-alignment, resulting in elevated BER across SNR levels. Notably, in the high-resolution setting, the proposed DP–SQ mechanism achieves BER performance nearly indistinguishable from the non-private baseline, demonstrating that strong privacy can be attained with minimal loss when sufficient angular resolution is available.
In Fig. 8, we visualize the received constellation points to provide intuition behind these trends. SVD-based beamforming produces tightly clustered decision points around the ideal 16-QAM constellation. Deterministic quantization introduces moderate angular distortion, broadening the symbol clouds. The DP–SQ mechanism injects controlled stochasticity into the angular representation, further dispersing the received symbols while offering formal DP guarantees.
@@
Next, in Fig. 9, we report the average beamforming gain $G$ over $5000$ Rayleigh fading realizations for several IEEE 802.11 antenna configurations with $(N_{t},N_{r})\in{(2,1),(2,2),(2,3),(2,4),(2,8)}$ and a single spatial stream ($N_{s}=1$). Further, Fig. 9 reports the mean gain achieved by three schemes: (i) deterministic mid-point quantization of the Givens angles, (ii) the proposed differentially private stochastic quantization (DP–SQ), and (iii) BeamDancer-style randomized beamforming [2]. The perfect SVD beamformer provides the reference gain $G=1$. Deterministic quantization incurs moderate loss due to finite angular resolution. The proposed DP–SQ mechanism introduces controlled stochastic perturbations governed by the privacy parameters $(\varepsilon_{\phi},\varepsilon_{\psi})$, producing a tunable privacy–utility tradeoff: smaller values of $\varepsilon$ increase privacy but yield additional beam mis-alignment and corresponding gain reduction. BeamDancer [2] exhibits the largest degradation, as its high-variance integer-bin perturbations cause substantial angular distortion. Across all antenna configurations, the trend is consistent: privacy-driven angular randomization directly impacts achievable array gain, with the effect magnified in larger arrays, yet the proposed DP-SQ scheme preserves compatibility with the 802.11 feedback format while offering significantly improved privacy–utility tradeoff.
TABLE II: Default WiFi channel and system simulation parameters.
Parameter Symbol Value
MIMO Configuration
Transmit antennas $N_{t}$ 2
Receive antennas $N_{r}$ 1
Spatial streams $N_{s}$ 1
Antenna spacing $d$ $0.5\lambda$
OFDM / Bandwidth
Carrier frequency $f_{c}$ 5.785 GHz
Channel bandwidth $B$ 20 MHz
Number of subcarriers $N_{sc}$ 256
Subcarrier spacing $\Delta f$ 78.125 kHz
Channel Model
Rician $K$-factor $K$ 5 dB
Number of multipath taps $L$ 20
Maximum excess delay $\tau_{\max}$ 4 samples
User direction of arrival $\theta_{0}$ $15^{\circ}$
Temporal Sampling
Snapshot interval $\Delta t$ 1 ms
Number of snapshots $T$ 5000
Observation duration $T\Delta t$ 5 s
Robustness to Adversarial Estimation. We further verify our ability to perturb the adversary’s access to the legitimate user information. Table II summarizes the simulation parameters. We model an IEEE 802.11ac MIMO-OFDM link [12] at $f_{c}=5.785$ GHz with a $20$ MHz bandwidth and $N_{sc}=256$ subcarriers. The channel follows a Rician fading model ($K=5$ dB) with $L=20$ multipath taps, where the dominant line-of-sight component carries the user’s Doppler signature. CSI snapshots are sampled every $1$ ms over $5$ seconds, providing sufficient temporal resolution to capture the micro-Doppler shifts induced by human motion as the user transitions through stationary, walking, jogging, and running activity zones. Fig. 10 provides an example demonstration of the effectiveness of our proposed privacy-based perturbations over time: despite the injected noise on the compressed CSI, the normalized per-stream beamforming gain generally remains close. Fig. 11 further quantifies this, where the median gain obtained from the distorted CSI is $0.89$ (compared to the deterministic median gain of $0.97$ and never drops below $0.52$. A closer look at the CSI, as shown in Fig. 12 reveal how this occurs. Under deterministic quantization, the heatmap exhibits fine-grained frequency-selective fading patterns and rapid temporal fluctuations that encode micro-Doppler signatures correlated with user speed. However, under DP-SQ, these patterns are altered but remain well-correlated with the deterministic quantization. Thus, the overall beamforming gain is expected to be similar to the deterministic-quantization equivalent, and the speed estimation is expected to be degraded due to changes in the ground-truth temporal CSI structure. However, under DP-GSQ, the patterns are visibly distorted, obscuring the speed-dependent micro-Doppler structure while preserving the broadband channel envelope needed for beamforming.
We further validate this through a Monte Carlo study comprising 1000 independent trials, each with a randomly generated speed profile drawn from realistic human-motion parameters as defined in Table  III. Fig. 13a and Fig. 13b report the average and median beamforming gain, respectively, as a function of the randomization probability $p$. With probability 1p1-p the angle is mapped to its nearest quantization level (deterministic quantization), and with probability $p$ it is replaced by a sample drawn uniformly from the $k$ adjacent candidate levels surrounding the true value. Setting $p=0$ defaults to the standard scalar quantization with no privacy noise, while $p=1$ yields maximum randomization (uniform over the $k$-neighborhood). The deterministic baseline ($p=0$) achieves an average gain of $0.9870$ and a median gain of $0.9900$, both near unity, confirming that standard IEEE 802.11 quantization introduces negligible beamforming loss. As $p$ increases, both metrics degrade monotonically: at moderate privacy ($p=0.3$) the average and median gains remain above $0.75$ and $0.90$, respectively, whereas at full randomization ($p=1$) they drop to approximately $0.54$ and $0.58$. Fig. 13c presents the average speed-classification error of a micro-Doppler adversary attempting to infer the user’s activity zone from the quantized CSI feedback. At the deterministic baseline the adversary mis-classifies only $19.0\%$ of time windows, but the error rises steeply with $p$: it exceeds $43\%$ at $p=0.2$, surpasses $50\%$ at $p=0.3$, and saturates near $73\%$ for $p\geq 0.9$, approaching the $75\%$ error rate of a uniform random guess over the four activity classes. The error bars ($\pm 1\sigma$ across trials) are tight throughout, confirming that these trends are statistically robust. Taken together, these results demonstrate that the DP-SQ mechanism offers a practical operating region around $p\in[0.1,\,0.3]$ where the adversary’s classification ability is substantially degraded (error $30\%$) while the beamforming gain remains above $75\%$ of the ideal value.
@@
This indicates that the DP mechanism preserves most of the communication utility. At the same time, the same perturbations substantially degrade the eavesdropper’s ability to track the user’s motion, leading to visibly biased and inaccurate velocity trajectories compared to the ground truth. This illustrates the key privacy–utility tradeoff achieved by our design: the beamformer retains near-optimal precoding performance from the CSI privatized feedback, while an adversary observing the perturbed CSI suffers a pronounced loss in speed-estimation accuracy from the reduced information leakage.
@@
It is worth emphasizing that the privacy guarantee of the proposed mechanism is per-shot: each CSI report is individually protected under $\varepsilon$-DP. In our setup, the adversary obtains one CSI measurement every $10^{-3}$ seconds over a total communication duration of $T=10$ seconds, resulting in 10,00010{,}000 DP-protected releases. By applying the advanced composition theorem of DP [3], the cumulative privacy loss grows sublinearly, yielding an overall leakage on the order of $\varepsilon_{\mathrm{total}}={O}(\sqrt{T}\,\cdot\max\{\varepsilon_{\psi},\varepsilon_{\phi}\})$, demonstrating that even over long communication intervals, the aggregated DP leakage remains well controlled.